OpenStudy (anonymous):

(5mn)(-m^2n^2)(-2n^5p)

OpenStudy (kropot72):

First multiply the terms 5mn and -m^2n^2 to give -5m^3n^3. Then multiply this result by -2n^5p:

OpenStudy (kropot72):

\[(-5m ^{3}n ^{3})\times (-2n ^{5}p)=?\]

OpenStudy (anonymous):

what's the anwer ?

OpenStudy (kropot72):

The answer is found from the following breakdown: \[(-5\times -2)\times (m ^{3})\times (n ^{3}\times n ^{5})\times p=?\]

OpenStudy (anonymous):

Solve .. 3x4(-3x^2 ? Thankyou .